Lynch Company manufactures and sells a single product. The following costs were incurred during the company’s first year of operations:

Variable costs per unit:
Manufacturing:
Direct materials $ 13
Direct labor $ 7
Variable manufacturing overhead $ 1
Variable selling and administrative $ 1
Fixed costs per year:
Fixed manufacturing overhead $ 288,000
Fixed selling and administrative $ 198,000

During the year, the company produced 24,000 units and sold 20,000 units. The selling price of the company’s product is $48 per unit.

Required:

1. Assume that the company uses absorption costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.

Prepare an income statement for the year. Assume that the company uses absorption costing.

Lynch Company
Absorption Costing Income Statement


2. Assume that the company uses variable costing:

a. Compute the unit product cost.

b. Prepare an income statement for the year.

Prepare an income statement for the year. Assume that the company uses variable costing

Lynch Company
Variable Costing Income Statement
Answer #1

Answer 1 (a):

Unit product cost using absorption costing = $33.00

Working:

Variable manufacturing costs per unit: Direct material Direct labor Variable manufacturing overhead Fixed manufacturing overh

Fixed manufacturing overhead per unit = 288000 / 24000 = $12.00

Answer 1 (b):

Absorption Costing Income Statement

Lynch Company Absorption Costing Income Statement For the Year Ended XXXXXXXXXXXX Sales $960,000 Cost of Goods Sold Beginning

Working:

Sales = 20000 * $48 = $960,000

Cost of Goods manufactured = 24000 * $33 = $792,000

It is first year of operation; hence beginning Finished goods inventory = $0

Ending finished goods inventory = (24000 - 20000) * $33 = $132,000

Variable selling and administrative = 20000 * $1 = $20,000

Answer 2 (a):

Unit product cost using Variable costing = $21.00

Working;

Variable manufacturing costs per unit: Direct material Direct labor Variable manufacturing overhead Unit product cost using V

Answer 2 (b):

Variable Costing Income Statement:

Lynch Company Variable Costing Income Statement For the Year Ended XXXXXXXXXXXX Sales $960,000 Less: Variable costs Cost of G

Working:

Cost of Goods manufactured = 20000 * $21 = $420,000